Heat supplied to raise steam in the boiler is 3000 KJ/kg while heat rejected by the system to
the cooling water in the condenser is 2000 KJ/kg. The turbine develops power of 1500 KW and
the work required to pump the condensate back into the boiler is 50 KW.
1
Expert's answer
2020-06-02T17:52:11-0400

Actual pump Work "W = Q_{S}-Q_{R} = 3000-2000 = 1000 \frac{KJ}{Kg}"